OpenStudy (anonymous):

If a can of ravioli with a diameter of 2 inches and a height of 4.5 inches holds 12 ounces of ravioli, how many ounces of ravioli would a can with the same height, but double the diameter, hold?

OpenStudy (anonymous):

Ok, let's take a look. First of all, we need to use the formula for the cilinder's formula: V= pi*r²*h So in the original cilinder, this is V= pi*1²*4.5 = 9/2 and that equals 12 ounces In the second cilinder, the volume is this: V = pi*2²*4.5 = 18 So, 18 is 4 times 9/2 and 9/2 equals 12, so 18 equals 12*4 = 48 There fits 48 ounces opf ravioli in the can